OpenStudy (anonymous):

Graph the function f(x) = (x + 3)3 by hand and describe the end behavior.

OpenStudy (anonymous):

this function looks exactly like \(\large y=x^3 \). do you know how to graph this and the end behaviour of this?

OpenStudy (anonymous):

wait, the function is f(x)=(x+3)^3 or f(x)=3*(x+3)??

OpenStudy (anonymous):

You are not going to get an answer if you dont respond to the people that are trying to help you

OpenStudy (anonymous):

I don't understand what they are asking

OpenStudy (anonymous):

It's f(x) = (x+3)^3

OpenStudy (anonymous):

ok, then watch what dpaInc said above

OpenStudy (anonymous):

do you know what they're asking as far as "end behaviour" is concerned?

OpenStudy (anonymous):

the function you have have the exact same graph, end behaviour as \(\large y=x^3\) except that your function is translate 3 units to the